# SIGMA: Sheaf-Informed Geometric Multi-Agent Pathfinding

### v1

@@ Abstract
The Multi-Agent Path Finding (MAPF) problem aims to determine the shortest and collision-free paths for multiple agents in a known, potentially obstacle-ridden environment. It is the core challenge for robotic deployments in large-scale logistics and transportation. Decentralized learning-based approaches have shown great potential for addressing the MAPF problems, offering more reactive and scalable solutions. However, existing learning-based MAPF methods usually rely on agents making decisions based on a limited field of view (FOV), resulting in short-sighted policies and inefficient cooperation in complex scenarios. There, a critical challenge is to achieve consensus on potential movements between agents based on limited observations and communications. To tackle this challenge, we introduce a new framework that applies sheaf theory to decentralized deep reinforcement learning, enabling agents to learn geometric cross-dependencies between each other through local consensus and utilize them for tightly cooperative decision-making. In particular, sheaf theory provides a mathematical proof of conditions for achieving global consensus through local observation. Inspired by this, we incorporate a neural network to approximately model the consensus in latent space based on sheaf theory and train it through self-supervised learning. During the task, in addition to normal features for MAPF as in previous works, each agent distributedly reasons about a learned consensus feature, leading to efficient cooperation on pathfinding and collision avoidance. As a result, our proposed method demonstrates significant improvements over state-of-the-art learning-based MAPF planners, especially in relatively large and complex scenarios, demonstrating its superiority over baselines in various simulations and real-world robot experiments.
@@ I INTRODUCTION
Recently, the MAPF community has started looking to Multi-Agent Reinforcement Learning (MARL) to generate fast and scalable solutions [5, 6, 7]. Moreover, MARL has gained significant traction in multi-robot systems, where agents collaborate in decentralized settings to achieve global objectives [8, 9, 10, 11, 12, 13]. These learning-based approaches rely on decentralized planning under partial observability (i.e., each agent only observes its nearby environment, usually 11$\times$11 grid world), reducing computational complexity and enabling the network to tackle large-scale scenarios effectively [14]. However, the solutions generated by these learning-based methods are usually suboptimal, since they restrict the information available to agents, hindering their ability to avoid local minima and perform delicate joint behaviors. To improve the performance of learning-based MAPF planners, recent methods tried to augment the information available to agents by incorporating expert paths, designing communication schemes, or providing global map encodings [15, 16, 17]. These methods show significant improvements over previous learning-based methods, but there is still a remarkable performance gap between learning-based solutions and centralized optimization-based solutions [18].
@@ III-B3 Reward
Move (Up/Down/Left/Right) $-0.075$
Stay (on goal, away goal) $0,-0.075$
Collision (obstacle/agents) $-0.5$
Finish $3$
@@ V-A Main Results Comparison
In our experiments, SIGMA consistently outperforms other learning-based planners in terms of SR across all tasks. Notably, as the number of agents increases, SIGMA’s improvement in SR significantly exceeds that of the baseline planners. For instance, in complex scenarios where most learning-based planners struggle or fail to solve the problems, such as with 128 agents on a 40$\times$40 map, SIGMA achieves a SR of 69%. Even more impressively, on a larger 60$\times$60 map, SIGMA’s SR reaches 80%. These results highlight the effectiveness of our approach where the agents, through achieving consensus, successfully avoid congestion and overcrowding, demonstrating robust performance even under challenging conditions. Additionally, it is noteworthy that on a smaller 20x20 map, although SIGMA exhibits a high SR, the AR isn’t as impressive. This indicates that while consensus effectively prevents congestion, it does not necessarily aid in navigating out of such congested scenarios efficiently.

### v2

@@ Abstract
The Multi-Agent Path Finding (MAPF) problem aims to determine the shortest and collision-free paths for multiple agents in a known, potentially obstacle-ridden environment. It is the core challenge for robotic deployments in large-scale logistics and transportation. Decentralized learning-based approaches have shown great potential for addressing the MAPF problems, offering more reactive and scalable solutions. However, existing learning-based MAPF methods usually rely on agents making decisions based on a limited field of view (FOV), resulting in short-sighted policies and inefficient cooperation in complex scenarios. There, a critical challenge is to achieve consensus on potential movements between agents based on limited observations and communications. To tackle this challenge, we introduce a new framework that applies sheaf theory to decentralized deep reinforcement learning, enabling agents to learn geometric cross-dependencies between each other through local consensus and utilize them for tightly cooperative decision-making. In particular, sheaf theory provides a mathematical proof of conditions for achieving global consensus through local observation. Inspired by this, we incorporate a neural network to approximately model the consensus in latent space based on sheaf theory and train it through self-supervised learning. During the task, in addition to normal features for MAPF as in previous works, each agent distributedly reasons about a learned consensus feature, leading to efficient cooperation on pathfinding and collision avoidance. As a result, our proposed method demonstrates significant improvements over state-of-the-art learning-based MAPF planners, especially in relatively large and complex scenarios, demonstrating its superiority over baselines in various simulations and real-world robot experiments. The code is available at https://github.com/marmotlab/SIGMA
@@ I INTRODUCTION
Recently, the MAPF community has started looking to Multi-Agent Reinforcement Learning (MARL) to generate fast and scalable solutions [5, 6, 7]. Moreover, MARL has gained significant traction in multi-robot systems, where agents collaborate in decentralized settings to achieve global objectives [8, 9, 10, 11, 12, 13]. These learning-based approaches rely on decentralized planning under partial observability (i.e., each agent only observes its nearby environment, usually $\times11$ grid world), reducing computational complexity and enabling the network to tackle large-scale scenarios effectively [14]. However, the solutions generated by these learning-based methods are usually suboptimal, since they restrict the information available to agents, hindering their ability to avoid local minima and perform delicate joint behaviors. To improve the performance of learning-based MAPF planners, recent methods tried to augment the information available to agents by incorporating expert paths, designing communication schemes, or providing global map encodings [15, 16, 17]. These methods show significant improvements over previous learning-based methods, but there is still a remarkable performance gap between learning-based solutions and centralized optimization-based solutions [18].
@@ III-B3 Reward
Move (Up/Down/Left/Right) 0.075-0.075
Stay (on goal, away goal) 0,0.0750,-0.075
Collision (obstacle/agents) 0.5-0.5
Finish 33
@@ V-A Main Results Comparison
In our experiments, SIGMA consistently outperforms other learning-based planners in terms of SR across all tasks. Notably, as the number of agents increases, SIGMA’s improvement in SR significantly exceeds that of the baseline planners. For instance, in complex scenarios where most learning-based planners struggle or fail to solve the problems, such as with 128 agents on a $\times40$ map, SIGMA achieves a SR of 69%. Even more impressively, on a larger $\times60$ map, SIGMA’s SR reaches 80%. These results highlight the effectiveness of our approach where the agents, through achieving consensus, successfully avoid congestion and overcrowding, demonstrating robust performance even under challenging conditions. Additionally, it is noteworthy that on a smaller 20x20 map, although SIGMA exhibits a high SR, the AR isn’t as impressive. This indicates that while consensus effectively prevents congestion, it does not necessarily aid in navigating out of such congested scenarios efficiently.